Suppose C is a 3 x 3 matrix such that det (C) = 4 . Find det (C +C)


1
Expert's answer
2020-09-21T16:19:33-0400

Let's consider C is a matrix given by

"\begin{bmatrix}
 a & b&c \\
 d&e&f\\
g&h&i
\end{bmatrix}"

them determinant of matrix C can be written as


"\begin{vmatrix}
 a & b&c \\
 d&e&f\\
g&h&i
\end{vmatrix}=4.........(1)"


Now,

"det(C+C)=\begin{vmatrix}
 a & b&c \\
 d&e&f\\
g&h&i
\end{vmatrix}+\begin{vmatrix}
 a & b&c \\
 d&e&f\\
g&h&i
\end{vmatrix}="


"=\begin{vmatrix}
 2a & 2b&2c \\
 2 d&2e&2f\\
2g&2h&2i
\end{vmatrix}=2\times2\times2\times\begin{vmatrix}
 a & b&c \\
 d&e&f\\
g&h&i
\end{vmatrix}="


"=8\times4\space" from eq.(1)"="

"=32"

Answer:

"det(C+C)=32"